Question 57: A sample of 200 users of Cannondale bicycles was surveyed to ask them

then whether they would describe themselves as physically fit, and whether or not they use their bicycles primarily on weekends. The table below contains the result.

Uses bicycle primarily on weekends Yes   No

Physically Fit                                     20    80

Not Physically Fit                               70    30

If the sample is a good representation of the population of Cannondale bicycle users, we can expect that ________ percent of the population of Cannondale users to be physically fit and use their bicycles primarily on weekend:

  1. 10
  2. 45
  3. 50
  4. 20
Answer #1

A. 10

From the table we can see that 20 user uses their bicycle on weekends and consider themselves to be physically fit. Let X Be the required percentage

X/100*200= 20

X= 20/200*100 = 10